# GoSurf: Identifying Software Supply Chain Attack Vectors in Go

### v1

@@ Abstract.
In Go, the widespread adoption of open-source software has led to a flourishing ecosystem of third-party dependencies, which are often integrated into critical systems. However, the reuse of dependencies introduces significant supply chain security risks, as a single compromised package can have cascading impacts. Existing supply chain attack taxonomies overlook language-specific features that can be exploited by attackers to hide malicious code. In this paper, we propose a novel taxonomy of 12 distinct attack vectors tailored for the Go language and its package lifecycle. Our taxonomy identifies patterns in which language-specific Go features, intended for benign purposes, can be misused to propagate malicious code stealthily through supply chains. Additionally, we introduce GoSurf, a static analysis tool that analyzes the attack surface of Go packages according to our proposed taxonomy. We evaluate GoSurf on a corpus of widely used, real-world Go packages. Our work provides preliminary insights for securing the open-source software supply chain within the Go ecosystem, allowing developers and security analysts to prioritize code audit efforts and uncover hidden malicious behaviors.
copyright: none
@@ 1. Introduction
Software development today heavily relies on open-source software (OSS) dependencies, leveraging the benefits of code reusability. These dependencies enable developers to re-use pre-written, tested, and community-vetted code, significantly reducing development time and effort. This phenomenon is observed in all software ecosystems, including Java, Python and Go.
@@
The Go programming language, known for its simplicity, concurrency, and performance, is not immune to supply chain attacks. Its supply chain demands specific attention due to its growing adoption in security-critical software like Kubernetes (Kubernetes, 2024) and Go Ethereum (Ethereum, 2024). Indeed, there have already been advanced persistent threats (APTs) reported to use Go (Cimpanu, 2021), such as the second-stage malware backdoor Sunshuttle (Lindsay Smith, 2021), connected to the 2020 SolarWinds supply chain attack (for Cyber Security Denmark, 2021) targeting US government agencies and private companies. Similarly, a Golang backdoor GateDoor was leveraged in a 2024 supply chain attack (HackerNews, 2024) affecting downstream users of a video recording service.
@@
  • A novel taxonomy with 12 code injection attack vectors specifically tailored for the Go software stack, showcased by runnable proof of concepts.
@@
  • An analysis of open-source, real-world Go packages to evaluate GoSurf. We publicly release GoSurf and the evaluation scripts  [^1] https://anonymous.4open.science/r/SCORED24_GoSurf/.
@@ 2. Background on Golang
The Go programming language (The Go Programming Language, 2024d), commonly known as Golang, was created by Google in 2007. It is a systems language that addresses shortcomings of other programming languages by blending their distinctive characteristics (Pike, 2012), such as the static typing and runtime efficiency of C, the readability and usability of Python, and the high-performance multiprocessing features of Erlang. From a security standpoint, Go offers features like garbage collection, bounds checking, and type safety, which help prevent common vulnerabilities such as buffer overflows and memory leaks. Since its public release, Go has gained widespread adoption (Stack Overflow, 2023).
In Go, software is organized into units called Go modules, which are collections of related Go packages that are versioned and distributed together (The Go Programming Language, 2024c). A Go package consists of one or more source files within the same directory. Projects can depend on entire modules or only on single packages, according to the specific needs of their application.
Open-source Go modules are supplied to dependents through a distributed package management system. By publishing a Go module to a version control system (VCS) such as GitHub, any developer can import it by referencing its repository hosting location (The Go Programming Language, 2024c). An official Go proxy (The Go Programming Language, 2024b) maintains mirrors for the repositories, ensuring that modules are reliably available in the Go package ecosystem.
@@ 1. Pre-Build
To create a Go module, developers can use the ’go mod init’ command, which generates a manifest file named ’go.mod’ (shown in Listing 1). This file lists all the required direct and indirect dependencies and ensures reproducible builds by pinning dependencies to specific versions (The Go Programming Language, 2024c). During this phase, dependencies can be explicitly integrated into the Go module through the ’go get’ command, which fetches and caches the required packages locally. After this, developers can test integrated packages (’go test’) and utilize code generation tools (’go generate’) to automate tasks like creating boilerplate code or generating environment-specific artifacts, before the build phase.
@@ 2. Build
During the compilation of a module, the ’go build’ tool parses the manifest (’go.mod’) and automatically fetches any required dependencies not cached locally. The Go compiler then compiles them into machine code (The Go Programming Language, 2024g). The compiled code is linked with the Go runtime to create an executable binary. Go toolchain is explicitly designed to avoid code execution during this process, ensuring build-time security.
@@ 3. Initialization
When executing the installed binary, the Go Runtime first initializes dependencies based on a dependency ordering (The Go Programming Language, 2024d). If package A imports package B, package B will be initialized before package A. First, package-level global variables are initialized, in the order they are declared. Then, the Go runtime executes any package-level init() functions. Global variables initialization and init functions can contain arbitrary code.
@@ 4. Execution
Finally, once all packages are initialized, the Go runtime will execute the ’main’ function in the main package, as a goroutine (The Go Programming Language, 2024d). A goroutine is a lightweight thread managed by the Go runtime, allowing for concurrent execution with lighter overhead than with traditional OS threads. The Go runtime also automatically manages memory allocation and garbage collection.
@@ 3. Supply Chain Attack Vectors in Go
Establishing a robust taxonomy of attack vectors is critical in the realm of software supply chain security. Although previous efforts (Ladisa et al., 2023a, b; Ohm et al., 2020) have yielded general taxonomies for supply chain attack strategies, they are not tailored to the unique characteristics and idiosyncrasies of specific programming languages and stacks. The Go programming language includes features and design choices that both mitigate and introduce distinct security challenges.
In this section, we propose a novel language-specific taxonomy for classifying attack vectors in the supply chain of Go packages. With this taxonomy, we can help researchers and security experts to effectively understand and identify the unique risks linked to Go dependencies.
To develop a Go-specific taxonomy of supply chain attack vectors, we began by thoroughly reviewing existing generic taxonomies (Ladisa et al., 2023a, b; Ohm et al., 2020) and systematically questioning the meaning and relevance of each attack vector in the context of Go. We then curated a dataset of runnable proof-of-concept attacks written in Go (included in the project repository) to assess their feasibility and associated risks.
@@ 3.1. Malicious Code at Pre-Build Time
Test suites create a low-visibility environment with potentially large amounts of code, allowing attackers to hide malicious code for seamless execution during build integration. This code might compromise either the developer machine or the CI runners.
Testing functions are often overlooked in security reviews because of the absence of business logic that ends up in production. However, the same testing functions may be triggered during post-deployment smoke tests. This provides attackers with the opportunity to execute arbitrary code on production machines.
@@ 3.3. Malicious Code at Execution Time
In Go, developers define custom functions that serve as constructors for structures following a naming convention such as ’NewStructName’ (The Go Programming Language, 2024e). If the variable being initialized is the primary type for the given package, the constructor can be named ’New’ without suffix.
@@
Constructors are called to create a new instance of a structure, and any malicious code embedded into constructors will execute during each instantiation. This exploit is insidious, as constructors can be assumed to perform benign and simple instantiation tasks, potentially leading to less developer scrutiny. This makes hiding malicious code in constructors an effective method for attackers to compromise a system repeatedly.
@@
Reflection in Go enables dynamic inspection and runtime modification of types and methods. This is accomplished through the reflect package. When a method is invoked through reflection, the actual method called is determined at runtime rather than at compile-time. Developers use reflection to build generic libraries, implement serialization and deserialization mechanisms, and create flexible APIs that can operate on various types without needing to know their specifics at compile-time (The Go Programming Language, 2024h).
@@
An attacker can exploit reflection to dynamically inject and execute arbitrary code at runtime. For instance, consider the scenario shown in Listing 4; reflection is used to call ’UnsafeMethod’ dynamically.
@@
The method name to be called is stored in a variable that an attacker can manipulate (lines 10-11) to execute arbitrary methods. This ability to invoke methods dynamically makes it easier for attackers to hide malicious payloads and evade static security analysis (Sayar et al., 2023).
@@ 2 Execute()
Safe usage of unsafe pointers is possible, however, several examples of patterns that introduce security risks exist (Costa et al., 2022; Lauinger et al., 2020). First, attackers can exploit unsafe pointers to perform unauthorized memory operations. As shown in Listing 6, unsafe pointers can be misused to create a function pointer variable and set its value to the address of an arbitrary function (lines 7-8), thereby effectively allowing the execution of any code at runtime. Similarly, Listing 7 shows how unsafe pointers can be misused to access out-of-bound memory locations, potentially leading to information disclosure or memory-corruption vulnerabilities.
@@
1func targetFunction() {
2 // malicious code
3}
4func main() {
@@
Go supports static linking of assembly code (The Go Programming Language, 2024f). Developers can write package-level assembly functions using Go Assembly. When compiling, any assembly file is assembled into object files by the Go assembler, and then linked into the Go objects. These functions can be used in Go files by defining a function stub with the same name as that specified in the assembly file. The assembly functions can then be called as any normal function in Go (shown in Listing 9). Assembly is commonly leveraged for code optimization or in security-critical domains such as cryptography (Pike, 2016).
@@ 4. GoSurf
[…]
@@ 4.1. Design
[…]
GoSurf tool architecture design
The parser component parses Go source code and constructs an Abstract Syntax Tree (AST). The static analysis engine uses the AST and a set of rules to identify attack vectors based on the defined taxonomy. The reporting component consolidates the analysis results presenting them in a user-friendly format.
We implemented GoSurf in Go, using 793 lines of code, and leveraging Go standard libraries. The details of each component are explained in the following.
@@ 4.2.1. Pre-integration Auditing
Before integrating a third-party dependency, developers should conduct a thorough security audit. GoSurf offers a comprehensive overview of the risks associated with a dependency by reporting on its attack surface. This helps dependency consumers make informed decisions about which dependencies to include or avoid. Additionally, GoSurf streamlines the security auditing process by highlighting riskier portions of the codebase, characterized by a higher density of attack vectors, in order to prioritize auditing efforts.
@@ 5.1. Attack Surface in Real-World Go Modules
To analyze the applicability of our taxonomy, we conducted a comprehensive analysis using GoSurf across 10 Go projects. We carefully selected these projects based on their number of dependents, indicating their relevance and popularity. The selected projects cover diverse categories such as cloud orchestration, networking, blockchain, monitoring, and CLI applications, sourced from libraries.io. This ensures that we focus on modules with widespread real-world usage, which if exploited, would have large security implications. To limit sampling bias issues we performed the same analysis on a sample of the top 500 most imported packages obtaining the same findings as described in the following. Due to space constraints, detailed results on the entire dataset are available in the project repository.
@@ Results
Using GoSurf, we identified occurrences of some of the 12 attack vectors in each of the chosen Go modules. These are summarized in Table 1. The results indicate that P2 (testing functions), E1 (constructors methods), and E3 (interfaces) are the most frequently occurring attack vectors. Although their high prevalence may be connected to common benign Go programming practices, their usage call for additional auditing in a security-sensitive context. Overall, Table 1 shows that malicious actors have ample room to choose where to hide malicious code execution, as argued in Sec. 3.
Table 1. Usage of Taxonomy Attack Vectors in Popular Go Projects
Go Module Dep. LOC P1 P2 I1 I2 E1 E2 E3 E4 E5 E6 E7 E8
terraform@v1.8.5 1,384 377,115 38 2,899 238 50 4,383 121 3,707 10 0 0 0 17
kubernetes@v1.30.2 2,024 3,728,835 119 10,339 13,941 1,108 38,813 1,528 85,277 8,105 803 1,495 1 230
go-ethereum@v1.14.5 3,965 235,019 34 2,132 1,116 76 6,470 128 6,343 26 21 12 0 22
coredns@v1.11.1 642 52,348 3 534 118 98 784 18 603 3 0 0 0 1
ginkgo@v2.19.0 15,010 40,768 0 90 283 7 447 12 734 23 2 0 0 24
testify @v1.7.1 90,337 14,084 4 426 24 0 225 7 277 6 0 0 0 1
vault@v1.17.0 1,235 436,387 42 3,132 362 55 4,351 207 5,683 0 0 0 0 14
logrus@v1.9.3 38,020 4,414 0 127 1 4 73 1 161 0 0 0 0 1
prometheus@v2.52.0 1,097 173,422 5 1,186 178 50 2,206 18 3,961 11 0 0 0 12
cobra@v1.8.1 33,747 11,861 0 271 1 1 12 2 277 0 0 0 0 2
go-textseg@v15.0.0 617 22,693 6 3 9 0 7 2 0 0 0 0 0 2
Furthermore, the use of cross-language features such as C code and assembly (E5 and E6) in projects like Kubernetes and Go-Ethereum suggests that developers may trade the built-in security mechanisms in Go for less secure code, growing the attack surface. Although less common, the use of plugins (E7) is also present in the Kubernetes project. Here, the loading and execution of externally pre-built code requires thorough security analysis.
@@
Now let us consider Listing 12, demonstrating a real-world example of a dynamic external execution attack vector (E8) discovered by GoSurf in a Kubernetes dependency mistifyio/go-zfs. This small library implements a wrapper for ZFS command line tools. It is a third-party dependency maintained by 23 common GitHub users with no affiliation to the Kubernetes project. A malicious actor could gain access and inject malicious code to subvert the arguments of the Run function (line 1), leading to arbitrary code execution. This potential supply chain attack would have a cascading impact on all the dependents of the go-zfs library, including Kubernetes.
For attackers, it is tempting to target widely-used dependencies to maximize the impact on downstream applications. Dependencies exhibiting a large attack surface may be easier targets for attackers due to their feasibility. Based on our sample of targets, we observe that the project size tends to correlate with the size of the attack surface. However, we also note that go-textseg exhibits a significantly smaller attack surface compared to testify, despite having a larger codebase. This suggests that Go modules can adhere closely to coding practices that minimize security risks defined by our taxonomy.

### v2

@@ Abstract.
In Go, the widespread adoption of open-source software has led to a flourishing ecosystem of third-party dependencies, which are often integrated into critical systems. However, the reuse of dependencies introduces significant supply chain security risks, as a single compromised package can have cascading impacts. Existing supply chain attack taxonomies overlook language-specific features that can be exploited by attackers to hide malicious code. In this paper, we propose a novel taxonomy of 12 distinct attack vectors tailored for the Go language and its package lifecycle. Our taxonomy identifies patterns in which language-specific Go features, intended for benign purposes, can be misused to propagate malicious code stealthily through supply chains. Additionally, we introduce GoSurf, a static analysis tool that analyzes the attack surface of Go packages according to our proposed taxonomy. We evaluate GoSurf on a corpus of 500 widely used, real-world Go packages. Our work provides preliminary insights for securing the open-source software supply chain within the Go ecosystem, allowing developers and security analysts to prioritize code audit efforts and uncover hidden malicious behaviors.
Open-Source Security, Supply Chain Attacks, Golang
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 Workshop on Software Supply Chain Offensive Research and Ecosystem Defenses ; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 Workshop on Software Supply Chain Offensive Research and Ecosystem Defenses (SCORED ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3689944.3696166isbn: 979-8-4007-1240-1/24/10ccs: Security and privacy Malware and its mitigation
@@ 1. Introduction
Software development today heavily relies on open-source software (OSS) dependencies, leveraging the benefits of code reusability. These dependencies enable developers to reuse pre-written, tested, and community-vetted code, reducing development time and effort. This phenomenon is observed in all software ecosystems, including Java, Python and Go.
@@
The Go programming language, known for its simplicity, concurrency, and performance, is not immune to supply chain attacks. Its supply chain demands specific attention due to its growing adoption in security-critical software like Kubernetes (Kubernetes, 2024) and Go Ethereum (Ethereum, 2024). Indeed, there have already been reports of typosquatted Go packages, such as the widely used ’cli’ package, compromised in a malicious fork that introduced a malicious ’init’ function to collect private system information and exfiltrate it to a remote host (Henriksen, 2021).
@@
  • A novel taxonomy with 12 attack vectors for code execution specifically tailored for the Go software stack, showcased by runnable proof of concepts.
@@
  • An analysis of open-source, real-world Go packages to evaluate GoSurf. We publicly release GoSurf and the evaluation scripts  [^1] https://github.com/chains-project/GoSurf.
@@ 2. Background on Golang
The Go programming language (The Go Programming Language, 2024c), commonly known as Golang, was created by Google in 2007. It is a systems language that addresses shortcomings of other programming languages by blending their distinctive characteristics (Pike, 2012), such as the static typing and runtime efficiency of C, the readability and usability of Python, and the high-performance multiprocessing features of Erlang. From a security standpoint, Go offers features like garbage collection, bounds checking, and type safety, which help prevent common vulnerabilities such as buffer overflows and memory leaks. Since its public release, Go has gained widespread adoption (Stack Overflow, 2023).
In Go, software is organized into units called Go modules, which are collections of related Go packages that are versioned and distributed together (The Go Programming Language, 2024b). A Go package consists of one or more source files within the same directory. Projects can depend on entire modules or only on single packages, according to the specific needs of their application.
Open-source Go modules are supplied to dependents through a distributed package management system. By publishing a Go module to a version control system (VCS) such as GitHub, any developer can import it by referencing its repository hosting location (The Go Programming Language, 2024b). An official Go proxy maintains mirrors for the repositories, ensuring that modules are reliably available in the Go package ecosystem.
@@ 1. Pre-Build
To create a Go module, developers can use the ’go mod init’ command, which generates a manifest file named ’go.mod’ (shown in Listing 1). This file lists all the required direct and indirect dependencies and ensures reproducible builds by pinning dependencies to specific versions (The Go Programming Language, 2024b). During this phase, dependencies can be explicitly integrated into the Go module through the ’go get’ command, which fetches and caches the required packages locally. After this, developers can test integrated packages (’go test’) and utilize code generation tools (’go generate’) to automate tasks like creating boilerplate code or generating environment-specific artifacts, before the build phase.
@@ 2. Build
During the compilation of a module, the ’go build’ tool parses the manifest (’go.mod’) and automatically fetches any required dependencies not cached locally. The Go compiler then compiles them into machine code (The Go Programming Language, 2024f). The compiled code is linked with the Go runtime to create an executable binary. Go toolchain is explicitly designed to avoid code execution during this process, ensuring build-time security.
@@ 3. Initialization
When executing the installed binary, the Go Runtime first initializes dependencies based on a dependency ordering (The Go Programming Language, 2024c). If package A imports package B, package B will be initialized before package A. First, package-level global variables are initialized, in the order they are declared. Then, the Go runtime executes any package-level init() functions. Global variables initialization and init functions can contain arbitrary code.
@@ 4. Execution
Finally, once all packages are initialized, the Go runtime will execute the ’main’ function in the main package, as a goroutine (The Go Programming Language, 2024c). A goroutine is a lightweight thread managed by the Go runtime, allowing for concurrent execution with lighter overhead than with traditional OS threads. The Go runtime also automatically manages memory allocation and garbage collection.
@@ 3. Supply Chain Attack Vectors in Go
A supply chain attack typically involves two main stages: first, an attacker introduces malicious code in an upstream software component, which subsequently is propagated to downstream users. The second stage is the execution of this malicious code. In this paper, we focus on supply chain attacks related to source code, and define ‘attack vector’ as follows.
Definition
An attack vector is a language feature that can be potentially used by an attacker to hide and execute malicious code during the application lifecycle, at some point between build and production.
Those source code attack vectors are highly programming-language specific. We note that the usage of these features is not inherently malicious, and they are indeed mostly used for good. Yet, we claim that as potential attack vectors, they should be prioritized for security audits, as they are more prone to hide of execution of malicious code. Establishing a robust taxonomy of these attack vectors is critical in software supply chain security. Although previous efforts (Ladisa et al., 2023b; Ohm et al., 2020) have yielded general taxonomies for supply chain attack strategies for arbitrary code execution, they are not tailored to the unique characteristics and idiosyncrasies of specific programming languages and stacks.
In this section, we propose a novel language-specific taxonomy to classify attack vectors for code execution in Go packages. With this taxonomy, we can help researchers and security experts to effectively understand and identify the unique risks linked to Go dependencies. To develop a Go-specific taxonomy of supply chain attack vectors, we began by collecting two existing Go-specific vectors from the literature (Ladisa et al., 2023b) (constructors and init functions). Next, we reviewed known attack vectors in other languages (Ohm et al., 2020) and their relevance in Go, identifying an additional applicable vector (testing functions). Finally, by analyzing the Go documentation, we identified nine other vectors relying on Go-specific features, not covered by any previous work. To validate these findings, we curated a dataset of executable proof-of-concept attacks written in Go, included in the project repository.
@@ 3.1. Malicious Code at Pre-Build Time
Test suites create a low-visibility environment with potentially large amounts of code, allowing attackers to hide malicious code for seamless execution during build integration. This code might compromise either the developer machine or the CI runners. Testing functions may be overlooked in security reviews because of the absence of business logic that ends up in production. However, some testing functions can be triggered during post-deployment smoke tests. This provides attackers with the opportunity to execute arbitrary code on production machines.
@@ 3.3. Malicious Code at Execution Time
In Go, developers define custom functions that serve as constructors for structures following a naming convention such as ’NewStructName’ (The Go Programming Language, 2024d). If the variable being initialized is the primary type for the given package, the constructor can be named ’New’ without suffix.
@@
Constructors are called to create a new instance of a structure, and any malicious code embedded into constructors will execute during each instantiation. Since constructors are often expected to only handle simple initialization tasks, they may be overlooked as potential vectors for malicious activity. This makes hiding malicious code in constructors an effective method for attackers to compromise a system repeatedly.
@@
Reflection in Go enables dynamic inspection and runtime modification of types and methods. This is accomplished through the reflect package. When a method is invoked through reflection, the actual method called is determined at runtime rather than at compile-time. Developers use reflection to build generic libraries, implement serialization and deserialization mechanisms, and create flexible APIs that can operate on various types without needing to know their specifics at compile-time (The Go Programming Language, 2024g).
@@
An attacker can exploit reflection to dynamically inject and execute arbitrary code at runtime. For instance, consider the scenario shown in Listing 4; reflection is used to call ’UnsafeMethod’ dynamically. The method name to be called is stored in a variable that an attacker can manipulate (lines 10-11) to execute arbitrary methods. This ability to invoke methods dynamically makes it easier for attackers to hide malicious payloads and evade static security analysis (Sayar et al., 2023).
@@ 2 Execute()
Safe usage of unsafe pointers is possible, however, several examples of patterns that introduce security risks exist (Costa et al., 2022; Lauinger et al., 2020). First, attackers can exploit unsafe pointers to perform unauthorized memory operations. As shown in Listing 6, unsafe pointers can be misused to create a function pointer variable and set its value to the address of an arbitrary function (lines 7-8), thereby effectively allowing the execution of any code at runtime.
Similarly, Listing 7 shows how unsafe pointers can be misused to access out-of-bound memory locations, potentially leading to information disclosure or memory corruption vulnerabilities.
@@
1func targetFunction() {// malicious code}
2func main() {
@@
Go supports static linking of assembly code (The Go Programming Language, 2024e). Developers can write package-level assembly functions using Go Assembly. When compiling, any assembly file is assembled into object files by the Go assembler, and then linked into the Go objects. These functions can be used in Go files by defining a function stub with the same name as that specified in the assembly file. The assembly functions can then be called as any normal function in Go (shown in Listing 9). Assembly is commonly leveraged for code optimization or in security-critical domains such as cryptography (Pike, 2016).
@@ 4. GoSurf
In this paper, we introduce GoSurf, a novel tool that analyzes the attack surface of Go modules according to the taxonomy of attack vectors described in Sec. 3. It reports the usage of language-specific features and programming idioms that can be leveraged for the corresponding attack vectors. GoSurf is meant to aid open-source dependency analysis and audit to find malicious code.
@@ 4.1. Design
The parser component parses Go source code and constructs an Abstract Syntax Tree (AST). The static analysis engine uses the AST and a set of rules to identify attack vectors based on the defined taxonomy. The reporting component consolidates the analysis results presenting them in a user-friendly format. We implemented GoSurf in Go, using 793 lines of code, and leveraging Go standard libraries. The details of each component are explained in the following.
@@ 4.2.1. Pre-integration Auditing
Before integrating a third-party dependency, developers should conduct a thorough security audit. GoSurf offers a comprehensive overview of the risks associated with a dependency and highlights those riskier areas of the codebase with a higher number of attack vectors. It also provides a valuable security metric for dependency consumers, helping them prioritize those dependencies with smaller attack surface.
@@ 5.1. Attack Surface in Real-World Go Modules
To analyze the applicability of our taxonomy, we conducted a comprehensive analysis using GoSurf across 500 Go modules. We selected these modules based on their number of dependents, specifically the top 500 most imported packages, indicating their relevance and popularity. This selection criteria ensures that we focus on modules with widespread real-world usage, which are natural targets of supply chain attacks. The selected projects cover diverse categories such as cloud orchestration, networking, blockchain, monitoring, and CLI applications, sourced from libraries.io.
@@ Results
Using GoSurf, we identified occurrences of some of the 12 attack vectors in each of the analyzed Go modules. Figure 3 reports the total number of occurrences for each vector across the 500 analyzed modules.
Figure 3. Go Attack Vector Prevalence in 500 Popular Go Projects. The rare ones are easier to forbid in order to secure a supply chain.
[…]
graph for Attack Vectors Occurrences on top 500 Go Package
The results indicate that P2 (testing functions), E1 (constructors methods), E3 (interfaces) and E4 (unsafe) are the most frequently occurring attack vectors. Although their high prevalence may be connected to common benign Go programming practices, their usage call for additional auditing in a security-sensitive context. Overall, Table 3 shows that malicious actors have ample room to choose where to hide malicious code execution, as argued in Sec. 3.
Furthermore, the widespread usage of cross-language features such as C code and assembly (E5 and E6) (also reported in security-critical projects like Kubernetes and Go-Ethereum) suggests that developers may trade the built-in security mechanisms in Go for less secure code, growing the attack surface. Although less common, the use of plugins (E7) is also present. Here, the loading and execution of externally pre-built code requires thorough security analysis.
Now, let us consider Listing 12, demonstrating a real-world example of a dynamic external execution attack vector (E8) discovered by GoSurf in a Kubernetes dependency mistifyio/go-zfs. This small library implements a wrapper for ZFS command line tools and is maintained by 23 common GitHub users with no affiliation to the Kubernetes project.
It is important to clarify that the reported example is not a command injection vulnerability exploitable through a user-controlled input, but rather a potential attack vector, according to the definition in Section 3. It would be a good place for attackers to hide code in order to perform a supply chain attack. For example, a malicious actor would compromise a repository (e.g., a pull request that introduces a hidden payload or alters code in a subtle way), altering the dependency control flow to taint the arguments passed to Run (line 1), enabling execution of custom and lower-level OS commands. Here, the use of exec.Command presents a risk of a supply chain compromise, potentially affecting all the dependents of the go-zfs library, including Kubernetes.
@@
For attackers, it is tempting to target widely-used dependencies to maximize the impact on downstream applications. In addition, dependencies with a larger attack surface are more likely to be targeted because they provide more opportunities for hiding the execution of malicious code. Therefore, these dependencies require closer scrutiny.